OpenStudy (anonymous):

PLease show me step by step how you would solve these three questions. Question 1: Add and then reduce to lowest terms. 7 3/5 + 4 1/10 Question 2: Divide and then reduce to lowest terms. 5/8 ÷ 3/4 Question 3: Sam wanted to leave an 18% tip on his $24 bill, what is the total amount that he has to pay?

OpenStudy (anonymous):

7 3/5 can be 7 6/10 so you add the numbers than the fractions 7+4=11 and 6/10 + 1/10 = 7/10 so your answer to number one is 11 7/10

OpenStudy (anonymous):

thx so much can u help with the others?

OpenStudy (anonymous):

im not 100% sure on number 2, but number 3 you can set up a proportion 18/100 = x/24 cross multiply to get 100x=432 divide both sides by 100 and get 4.32 and add that onto 24 so the answer is 28.32 dollas

OpenStudy (ghostgate):

For number 2, 5/8 divided by 3/4, we would first have to make the denominators equal one another. To do this we must multiply both numbers by a particular number until they both equal one another. For instance, 4 can go into 8 twice, which means we don't need to change 5/8, but due to how we multiplied 4 by 2 we must now multiply 3 by 2 as well. Which equals 6/8 or 1/3, but we will stick with 6/8. Now for dividing, first off we have 5/8 and 6/8, flip one or the other around and we'll get: 8/5 divided by 6/8, now multiply which should equal 48/40 or 1 8/40, now simplifying we would take it down to 1 1/5. I hope that helps!